10) The ionic product [HIOH], which is the equilibrium constant for the dissociation of water, is 1.00 101 mo/L at 25°C and 1.45 101 moP/I2 at for the process. b) Calculate the value of the ionic product at body temperature.
Answer #1

since deltaG= -RTlnK

and deltaG= deltaH- T*deltaS = -RT lnK

lnK= -deltaH/RT + deltaS/R

where K = ioninc prodict of water and T is inK

when the above equation is written for two different temperatures

lnK1= -deltaH/RT1+ deltaS/ R                         (1)

lnK2=- deltaH/RT2+ deltaS/R                           (2)

lnK2- lnK1= deltaH/r*(1/T1-1/T2)

ln(1.45/1)= deltaH/R*(1/298-1/303)

deltaH/R= 6710 and deltaH= 6710*8.314 J/mole. =55787 J/mole= 55.787 Kj/mole

substituting this value in Eq.1

ln(1*10-14)= -6710/298+deltaS/R

deltaS= 80.8 J/mole.K

at body temperature of 37 deg.c= 37+273= 310K

lnK= -6710/310+ 80.8/8.314

K= 6*10-6 mole2/L2